a rock falls from a tower that is 320 feet high. As it is falling, its height is given by the formula h=320-16t^2. How many seconds (in tenths) will it take to get to the ground.(h=0)?
My solution
0=-16t^2+320t
0=-4t(4t-80)
4t=0 or 4t-80=0
t=0 or t=20
0 doesn't answer the question so I get 20 seconds. The answer key on my study sheet says 4.5 seconds. I am not sure what I did wrong. Thanks!

Answer by Alan3354(69443)

a rock falls from a tower that is 320 feet high. As it is falling, its height is given by the formula h=320-16t^2. How many seconds (in tenths) will it take to get to the ground.(h=0)?
My solution
0=-16t^2+320t
0=-4t(4t-80)
4t=0 or 4t-80=0
t=0 or t=20
0 doesn't answer the question so I get 20 seconds. The answer key on my study sheet says 4.5 seconds.
--------------
0=-16t^2+320 Not 320 t
-----------
16t^2 = 320
t^2 = 20
t = sqrt(20) =~ 4.472 seconds